You are building a walkway of uniform width around a 100 foot by 60 foot swimming pool. After completing the walkway, you want to put a fence along the outer edge of the walkway. You have 450 feet of fencing to enclose the walkway. What is the maximum width of walkway?
Answer by edjones(8007)

2L+2w=P
2(100+x)+2(60+x)=450
200+2x+120+2x=450
4x+320=450
4x=130
x=32.5' maximum width of walkway.
.
Ed